# Pasadena, CA 91125, USA

### v1

@@ 4.1.2 $A_{4}$
The corresponding 3d $\mathcal{N}=2$ theory can be read off from the Schur index as before, which is given by the $U(1)^{4}$ ACSM theory. We now search for a superpotential deformation (given by dressed monopole operators) that breaks all but $U(1)_{A}$ symmetry and triggers RG flow to $\mathcal{N}=4$ SCFT in the IR.
@@
Table 2: Monopole operators in the 3d theory for $(A_{1},A_{4})$. Each row indicates the dressed monopole operator with magnetic charge $m$, dressed by each chirals by the power in the corresponding column. Operators included in the superpotential is indicated with a check mark.
There are 12 candidate dressed monopole operators that can appear in the superpotential, listed in Table 2. The global symmetry without any superpotential is $U(1)^{8}$, and we know there is no extra flavor symmetry in the IR. There are 22 possible candidates for $U(1)_{A}$ left unbroken by turning on 7 monopole deformations. By requiring the Coulomb and Higgs limits of the superconformal index to be trivial, we are left with only two cases (up to the symmetry $U(1)_{1}\leftrightarrow U(1)_{4},U(1)_{2}\leftrightarrow U(1)_{3}$ and $\Phi\leftrightarrow\widetilde{\Phi}$):
$U(1)_{\Phi_{1}}+U(1)_{\Phi_{3}},\quad U(1)_{\Phi_{1}}+U(1)_{\widetilde{\Phi}_{1}}-U(1)_{\Phi_{3}}.$ (65)
Figure 4: The absolute value of three sphere partition function for the 3d theory of $(A_{1},A_{4})$ is under varying mixing parameter $\nu$ with $U(1)_{A}=-U(1)_{\Phi_{1}}+U(1)_{\Phi_{3}}$. The function is minimized at $\nu=0$.
Among them, $F$ is maximized at $\nu=0$ only for the first choice (Fig. 4).
@@
by mixing with the gauge symmetry $U(1)_{3}$.
This is also supported by considering various limits of the superconformal index
$\begin{split}\mathcal{I}_{SCI}(\mathfrak{q},T)&=1-\mathfrak{q}+\left(T^{-1}+T\right)\mathfrak{q}^{\frac{3}{2}}-2\mathfrak{q}^{2}+T^{-1}\mathfrak{q}^{\frac{5}{2}}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q})&=1,\\ \mathcal{I}_{H}(T)&=1.\end{split}$ (68)
The superconformal index contains an extra SUSY multiplet contribution $(T+T^{-1})\mathfrak{q}^{\frac{3}{2}}$. The Coulomb and Higgs limits imply a trivial Coulomb and a Higgs branch.
The $T$-refined half index reproduces the conjectural Macdonald index of $(A_{1},A_{4})$ AD theory
@@
It agrees with the known expressions for the Macdonald index.
@@ 4.1.3 General Form
Based on these calculations, we conjecture that the superpotential deformation for general $n$ is
$\sum_{i=1}^{n-1}V_{2i}\Phi_{2i-1}\Phi_{2i+1}+\sum_{i=1}^{n}V_{2i}\widetilde{\Phi}_{2i-1}\widetilde{\Phi}_{2i+1}\\ +\sum_{i=1}^{n}V_{2i-1}\Phi_{2i-2}\widetilde{\Phi}_{2i}+\sum_{i=2}^{n}V_{2i-1}\widetilde{\Phi}_{2i-2}\widetilde{\Phi}_{2i}\\ +V_{2n-1}\widetilde{\Phi}_{2n-2}\Phi_{2n}$ (71)
The axial symmetry that survives this deformation is
@@
where $U(1)_{T_{2i}}$’s are the topological symmetries for the $2i$-th gauge node.
@@ 4.2 $(A_{1},A_{2n+1})$ theory
The BPS quiver of $(A_{1},A_{2n+1})$ AD theory is given by Figure 5. This theory has an $n$-dimensional Coulomb branch, and its Higgs branch is given by $\mathbb{C}^{2}/\mathbb{Z}_{n+1}$ with flavor symmetry given by $U(1)$ (for $n\geq 2$). Here we focus on $n\geq 2$ and the $(A_{1},A_{3})=(A_{1},D_{3})$ case will be dealt separately when we discuss $(A_{1},D_{2n+1})$ theory.
@@
The theory has $U(1)$ flavor symmetry, which corresponds to
@@
We find that this formula is identical to that of the half index of the A-twisted 3d $U(1)^{2n+2}$ Chern-Simons theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$s and $R$ symmetry are $\rho=\{2,2,\cdots,2,0\}$ for odd $n$ and $\rho=\{1,2,2,\cdots,2,1\}$ for even $n$.
The theory has a flavor symmetry
@@
Also, we have $4n+2$ chirals multiplets, all with trial $R$ charge 0. Their gauge charges are
@@ 4.2.1 $A_{5}$
from which we can extract $U(1)^{6}$ gauge theory description as described above.
@@
There are 12 dressed monopoles with $R=2$ neutral under $U(1)_{f}$, listed in Table 3. The global symmetry before turning on the superpotential is $U(1)^{10}$, and the superpotential should break all symmetry except for the $U(1)_{f}$ flavor symmetry in addition to $U(1)_{A}\subset SO(4)_{R}$ symmetry in the IR. By turning on dressed monopole operators in the superpotential as is given in Table 3, we preserve the following symmetry:
@@
The superconformal index and its Coulomb/Higgs limits are
$\begin{split}\mathcal{I}_{SCI}(\mathfrak{q},T,z)&=1-T\mathfrak{q}^{\frac{1}{2}}+T^{\frac{3}{2}}(z^{-1}+z)\mathfrak{q}^{\frac{3}{4}}+(-2+T^{2})\mathfrak{q}\\ &\qquad+(-T^{\frac{1}{2}}+T^{\frac{5}{2}})(z^{-1}+z)\mathfrak{q}^{\frac{5}{4}}+(T^{-1}+T^{3}(z^{-1}+1+z^{2}))\mathfrak{q}^{\frac{3}{2}}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q},z)&=1,\\ \mathcal{I}_{H}(T,z)&=1+T+(z^{-1}+z)T^{\frac{3}{2}}+T^{2}+(z^{-1}+z)T^{\frac{5}{2}}+(z^{-2}+1+z^{2})T^{3}+\cdots,\end{split}$ (85)
The superconformal index contains $T\mathfrak{q}^{\frac{1}{2}}$ which can be thought of as the contribution from the $U(1)_{f}$ moment map operator. The Higgs limit coincides with the Hilbert series of $\mathbb{C}^{2}/\mathbb{Z}_{3}$ given as
$\mathrm{HS}(\mathbb{C}^{2}/\mathbb{Z}_{3};z,T)=\frac{1-T^{3}}{(1-T)(1-zT^{\frac{3}{2}})(1-z^{-1}T^{\frac{3}{2}})}\ ,$ (86)
and we indeed find agreement. In fact, among all possible deformations, we find that our choice presented above is the only one that reproduces the correct Higgs branch Hilbert series.
Now, we can calculate the $T$-refined half index and find that it matches the proposed Macdonald index of the $(A_{1},A_{5})$ AD theory
@@ 4.2.2 $A_{7}$
From this, we can read off the corresponding 3d theory as before, as $U(1)^{8}$ ACSM theory. This gauge theory has a rather high rank, which makes the direct computation of the superconformal index infeasible. Instead, we give an indirect check via computing the half index under a suitable superpotential deformation and matching with the Macdonald index.
We turn on the following superpotential:
@@
We conjecture that this theory flows to an $\mathcal{N}=4$ SCFT in the IR with a trivial Coulomb branch and the Higgs branch given as $\mathbb{C}^{2}/\mathbb{Z}_{4}$, which corresponds to the $(A_{1},A_{7})$ theory. Here $V_{i}$ denotes a bare monopole charged +1+1 under $U(1)_{T_{i}}$, and $V_{i,j}$ denotes the one charged under both $U(1)_{T_{i}}$ and $U(1)_{T_{j}}$ by +1+1. We justify this by reproducing the Macdonald index from the half-index by identifying the flavor and $U(1)_{A}$ symmetry as
@@
Computing the $T$-refined half index, we get
@@
$\begin{split}I\!\!I\!&=1+qT+q^{2}\left(T^{2}\left(z+\frac{1}{z}+1\right)+2T\right)\\ &\quad+q^{3}\left(T^{3}\left(z+\frac{1}{z}+1\right)+T^{2}\left(z+\frac{1}{z}+2\right)+2T\right)\\ &\quad+q^{4}\left(T^{4}\left(z^{2}+\frac{1}{z^{2}}+z+\frac{1}{z}+1\right)+T^{3}\left(2z+\frac{2}{z}+2\right)+T^{2}\left(z+\frac{1}{z}+5\right)+2T\right)+\cdots,\end{split}$ (93)
agrees with the known expression for the Macdonald index.
@@ 4.2.3 General Form
$U(1)_{A}=\frac{1}{2}(U(1)_{T_{1}}+U(1)_{T_{2n+2}})+\sum_{i=1}^{n}U(1)_{T_{2i+1}}-\sum_{j=1}^{n-1}(U(1)_{\Phi_{2i}}-U(1)_{\widetilde{\Phi}_{2i+2}})\quad(\text{$n$ even})$ (95)
@@ 4.3 $(A_{1},D_{2n+1})$ theory
The $(A_{1},D_{2n+1})$ theory has an $n$-dimensional Coulomb branch and a Higgs branch given by $\mathbb{C}^{2}/\mathbb{Z}_{2}$. It has $SU(2)$ flavor symmetry, and the associated VOA is given by $\widehat{\mathfrak{su}(2)}_{-\frac{4n}{2n+1}}$. The BPS quiver for this theory is $G=D_{2n+1}$ given as Figure 6.
@@
The $SU(2)$ flavor symmetry in the BPS quiver corresponds to
@@
We have $4n+2$ chirals multiplets, all with $R$-charge 0. Their gauge charges are given as follows:
@@ 4.3.1 $D_{3}$
There are total $6$ monopoles with $R=2$ preserving $U(1)_{f}$, listed in Table 4. The global symmetry before turning on superpotential is $U(1)^{6}$. We turn on the superpotential given in Table 4 to break the global symmetry down to $U(1)_{f}\times U(1)_{A}$.
Then, we can single out the combination(up to symmetry $U(1)_{2}\leftrightarrow U(1)_{3}$ and $\Phi_{1}\leftrightarrow\widetilde{\Phi}_{1}$)
@@
with the superpotential consists of monopole operators in Table 4. Upon mixing with gauge symmetry, the axial charge can also be represented as
@@
In the superconformal index, $\chi^{SU(2)}_{\mathbf{3}}\mathfrak{q}^{\frac{1}{2}}$ comes from the $SU(2)_{f}$ moment map and extra $\mathfrak{q}$ corresponds to the $\mathcal{N}=4$ stress-energy tensor. The Coulomb and Higgs limits indeed reproduce the Hilbert series of the trivial Coulomb branch and Higgs branch of $\mathbb{C}^{2}/\mathbb{Z}_{2}$, which is given as
@@
We also evaluate the three-sphere partition function as a function of mixing parameter $\nu$ and indeed find that it is minimized at $\nu=0$, see Fig. 7. This shows that our identification of superconformal $R$ symmetry and $U(1)_{A}$ symmetry is indeed correct.
Figure 7: The absolute value of three sphere partition function for the 3d theory for $(A_{1},D_{3})$ as a function of mixing parameter $\nu$ with $U(1)_{A}=U(1)_{\widetilde{\Phi}_{1}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$
The $T$-refined half index with (108) is given as
$\begin{split}I\!\!I\!(q,T,z)&=\frac{1}{(q)_{\infty}^{4}}\sum_{m\in\mathbb{Z}^{4}}q^{m_{1}(m_{2}+m_{3})}q^{m_{1}+m_{2}+m_{3}}z^{m_{4}-m_{2}}T^{m_{2}+m_{3}}\\ &\quad\times\left(\prod_{i=1}^{4}(q^{1+m_{i}})_{\infty}\right)(q^{1+m_{1}}T)_{\infty}(q^{1+m_{2}+m_{3}-m_{4}})_{\infty}\\ &=(q)_{\infty}(qT)_{\infty}\sum_{m\in\mathbb{Z}^{4}}\frac{q^{m_{1}(m_{2}+m_{3})}q^{m_{1}+m_{2}+m_{3}}z^{m_{4}-m_{2}}T^{m_{2}+m_{3}}}{\left(\prod_{i=1}^{4}(q)_{m_{i}}\right)(qT)_{m_{1}}(q)_{m_{2}+m_{3}-m_{4}}}\ ,\end{split}$ (111)
@@
agrees with the previous results.

### v2

@@ 4.1.2 $A_{4}$
The corresponding 3d $\mathcal{N}=2$ theory can be read off from the Schur index as before; it is given by a $U(1)^{4}$ ACSM theory with eight chiral multiplets. We now search for a superpotential deformation given by dressed monopole operators that breaks all the $U(1)^{8}$ flavor symmetries except for $U(1)_{A}$ and triggers an RG flow to the $\mathcal{N}=4$ SCFT in the IR.
@@
Table 2: Monopole operators in the 3d theory for $(A_{1},A_{4})$. Each row indicates the dressed monopole operator with magnetic charge $m$, dressed by each chiral multiplets by the power in the corresponding column. Operators included in the superpotential is indicated with a check mark.
There are 12 candidate gauge invariant half-BPS dressed monopole operators which can appear in the superpotential, listed in Table 2. There are 22 possible candidates for $U(1)_{A}$ left unbroken by turning on 7 independent monopole operators. By requiring that the Coulomb and Higgs limits of the superconformal index are trivial, we are left with only two cases (up to the symmetry $U(1)_{1}\leftrightarrow U(1)_{4},U(1)_{2}\leftrightarrow U(1)_{3}$ and $\Phi\leftrightarrow\widetilde{\Phi}$):
$-U(1)_{\Phi_{1}}+U(1)_{\Phi_{3}},\quad U(1)_{\Phi_{1}}+U(1)_{\widetilde{\Phi}_{1}}-U(1)_{\Phi_{3}}.$ (64)
Figure 4: The absolute value of three-sphere partition function for the 3d theory of $(A_{1},A_{4})$ as a function of the mixing parameter $\nu$ with $U(1)_{A}=-U(1)_{\Phi_{1}}+U(1)_{\Phi_{3}}$. The function is minimized at $\nu=0$.
Among them, only for the first choice, we find that the $F$ is maximized at $\nu=0$ (see Figure 4) and the superconformal index at the fixed point gives
$\mathcal{I}_{SCI}(\mathfrak{q},T)=1-\mathfrak{q}+\left(T^{-1}+T\right)\mathfrak{q}^{\frac{3}{2}}-2\mathfrak{q}^{2}+T^{-1}\mathfrak{q}^{\frac{5}{2}}+\cdots.$ (65)
which agrees with the expectation that the supersymmetry is enhanced to $\mathcal{N}=4$ at the fixed point. One can also check that
$\mathcal{I}_{C}(\mathfrak{q})=1,\qquad\mathcal{I}_{H}(T)=1.$ (66)
as expected.
@@
by mixing with the gauge symmetry for the third gauge node, $U(1)_{3}$.
The $T$-refined half-index in the A-twisted limit reproduces the conjectural Macdonald index of $(A_{1},A_{4})$ AD theory
@@
It agrees with the known $q$-expansion of the Macdonald index.
@@ 4.1.3 General Form
Based on these calculations, we conjecture that the superpotential deformation for general $n$ as
$\begin{split}W&=\sum_{i=1}^{n-1}V_{2i}\Phi_{2i-1}\Phi_{2i+1}+\sum_{i=1}^{n}V_{2i}\widetilde{\Phi}_{2i-1}\widetilde{\Phi}_{2i+1}\\ &\quad+\sum_{i=1}^{n}V_{2i-1}\Phi_{2i-2}\widetilde{\Phi}_{2i}+\sum_{i=2}^{n}V_{2i-1}\widetilde{\Phi}_{2i-2}\widetilde{\Phi}_{2i}+V_{2n-1}\widetilde{\Phi}_{2n-2}\Phi_{2n}\ .\end{split}$ (71)
The axial symmetry survives under this deformation is
@@
where $U(1)_{T_{2i}}$ denotes the topological symmetries for the $2i$-th gauge node.
@@ 4.2 $(A_{1},A_{2n+1})$ theory
The BPS quiver of the $(A_{1},A_{2n+1})$ AD theory is given in Figure 5. This theory has an $n$-dimensional Coulomb branch, and its Higgs branch is given by $\mathbb{C}^{2}/\mathbb{Z}_{n+1}$ with $U(1)$ flavor symmetry (for $n\geq 2$). Here we focus on $n\geq 2$ and the $(A_{1},A_{3})=(A_{1},D_{3})$ case will be dealt separately when we discuss $(A_{1},D_{2n+1})$ theory.
@@
The $U(1)$ flavor symmetry corresponds to the flavor charge
@@
As before, we start from the observation is that this formula coincides with the half-index of a 3d $U(1)^{2n+2}$ Chern-Simons theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$s and $R$-symmetry are $\rho=\{2,2,\cdots,2,0\}$ for odd $n$ and $\rho=\{1,2,2,\cdots,2,1\}$ for even $n$.
The Higgs branch flavor symmetry can be idenfied with
@@
Also, we have $4n+2$ chiral multiplets, all with trial $R$-charge 0. Their gauge charges are
@@ 4.2.1 $A_{5}$
from which we can extract the $U(1)^{6}$ gauge theory description as described above.
@@
There are 12 dressed monopoles with $R=2$ that are neutral under $U(1)_{f}$, listed in Table 3. The global symmetry before turning on the superpotential is $U(1)^{10}$, and we expect that there exists a superpotential deformation that breaks all the global symmetries except for the $U(1)_{f}$ and the $U(1)_{A}$ symmetry in the IR. Among all possible deformations, we find that turning on the monopole operators indicated by checkmark in Table 3 preserves the following symmetry:
@@
The $F$-maximization is not feasible for this example, but one can start from the assumption that the Schur index (82), when interpreted as the half-index of the 3d $\mathcal{N}=4$ theory, is computed in the A-twisted limit and reverse engineer it to find the superconformal R-charge. The superconformal index computed with these data is
$\begin{split}\mathcal{I}_{SCI}(\mathfrak{q},T,z)&=1-T\mathfrak{q}^{\frac{1}{2}}+T^{\frac{3}{2}}(z^{-1}+z)\mathfrak{q}^{\frac{3}{4}}+(-2+T^{2})\mathfrak{q}\\ &\qquad+(-T^{\frac{1}{2}}+T^{\frac{5}{2}})(z^{-1}+z)\mathfrak{q}^{\frac{5}{4}}+(T^{-1}+T^{3}(z^{-1}+1+z^{2}))\mathfrak{q}^{\frac{3}{2}}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q},z)&=1,\\ \mathcal{I}_{H}(T,z)&=1+T+(z^{-1}+z)T^{\frac{3}{2}}+T^{2}+(z^{-1}+z)T^{\frac{5}{2}}+(z^{-2}+1+z^{2})T^{3}+\cdots.\end{split}$ (85)
The superconformal index contains the term $T\mathfrak{q}^{\frac{1}{2}}$, which can be thought of as the contribution from the $U(1)_{f}$ moment map operator. The Higgs limit coincides with the Hilbert series of $\mathbb{C}^{2}/\mathbb{Z}_{3}$ given as
$\mathrm{HS}(\mathbb{C}^{2}/\mathbb{Z}_{3};z,T)=\frac{1-T^{3}}{(1-T)(1-zT^{\frac{3}{2}})(1-z^{-1}T^{\frac{3}{2}})}\ .$ (86)
In fact, among all possible deformations, we find that our choice presented above is the only one that reproduces the correct Higgs branch Hilbert series.
Now, we can calculate the $T$-refined half-index in the A-twisted limit and find that it matches the proposed Macdonald index of the $(A_{1},A_{5})$ AD theory
@@ 4.2.2 $A_{7}$
From this, we can read off the corresponding 3d theory as before: a $U(1)^{8}$ ACSM theory with 14 chiral multiplets. This gauge theory has a rather high rank, which makes the direct computation of the superconformal index infeasible. Instead, as a non-trivial consistency check, we show that there exists a suitable superpotential deformation and a corresponding $U(1)_{A}$ symmetry that reproduces the conjecutural form of the Macdonald index.
We conjecture that there exists the following superpotential deformation
@@
and the theory flows to an $\mathcal{N}=4$ SCFT in the IR with a trivial Coulomb branch and the Higgs branch given as $\mathbb{C}^{2}/\mathbb{Z}_{4}$. Here $V_{i}$ denotes a bare monopole charged +1+1 under $U(1)_{T_{i}}$, and $V_{i,j}$ denotes the one charged under both $U(1)_{T_{i}}$ and $U(1)_{T_{j}}$ by +1+1. We justify this by reproducing the Macdonald index from the half-index by identifying the flavor and $U(1)_{A}$ symmetry as
@@
Computing the $T$-refined half-index in the A-twisted limit, we get
@@
$\begin{split}I\!\!I\!&=1+qT+q^{2}\left(T^{2}\left(z+\frac{1}{z}+1\right)+2T\right)\\ &\quad+q^{3}\left(T^{3}\left(z+\frac{1}{z}+1\right)+T^{2}\left(z+\frac{1}{z}+2\right)+2T\right)\\ &\quad+q^{4}\left(T^{4}\left(z^{2}+\frac{1}{z^{2}}+z+\frac{1}{z}+1\right)+T^{3}\left(2z+\frac{2}{z}+2\right)+T^{2}\left(z+\frac{1}{z}+5\right)+2T\right)+\cdots.\end{split}$ (93)
This agrees with the known expression for the Macdonald index.
@@ 4.2.3 General Form
$U(1)_{A}=\frac{1}{2}(U(1)_{T_{1}}+U(1)_{T_{2n+2}})+\sum_{i=1}^{n}U(1)_{T_{2i+1}}-\sum_{j=1}^{n-1}(U(1)_{\Phi_{2i}}-U(1)_{\widetilde{\Phi}_{2i+2}})\quad(\text{$n$ even})\ ,$ (95)
@@ 4.3 $(A_{1},D_{2n+1})$ theory
The $(A_{1},D_{2n+1})$ theory has an $n$-dimensional Coulomb branch and a Higgs branch given by $\mathbb{C}^{2}/\mathbb{Z}_{2}$. It has $SU(2)$ flavor symmetry, and the associated VOA is given by $\widehat{\mathfrak{su}(2)}_{-\frac{4n}{2n+1}}$. The BPS quiver for this theory is given in Figure 6.
@@
The $SU(2)$ flavor symmetry corresponds to
@@
We have $4n+2$ chiral multiplets, all with $R$-charge 0. Their gauge charges are given as follows:
@@ 4.3.1 $D_{3}$
There are $6$ monopoles with $R=2$ and preserving $U(1)_{f}$, listed in Table 4. The global symmetry before turning on the superpotential is $U(1)^{6}$. We turn on the superpotential given in Table 4 to break the global symmetry down to $U(1)_{f}\times U(1)_{A}$.
Then, we can single out the combinations (up to symmetry $U(1)_{2}\leftrightarrow U(1)_{3}$ and $\Phi_{1}\leftrightarrow\widetilde{\Phi}_{1}$)
@@
with the superpotential consisting of monopole operators in Table 4. Upon mixing with gauge symmetry, the axial charge can also be represented as
@@
In the superconformal index, $\chi^{SU(2)}_{\mathbf{3}}\mathfrak{q}^{\frac{1}{2}}$ comes from the $SU(2)_{f}$ moment map and the extra $\mathfrak{q}$ corresponds to the $\mathcal{N}=4$ stress-energy tensor. The Coulomb and Higgs limits indeed reproduce the Hilbert series of the trivial Coulomb branch and Higgs branch of $\mathbb{C}^{2}/\mathbb{Z}_{2}$, which is given as
@@
We also evaluate the three-sphere partition function as a function of the mixing parameter $\nu$ and indeed find that it is minimized at $\nu=0$, see Figure 7. This shows that our identification of superconformal $R$-symmetry and $U(1)_{A}$ symmetry is indeed correct.
Figure 7: The absolute value of three-sphere partition function for the 3d theory for $(A_{1},D_{3})$ as a function of the mixing parameter $\nu$ with $U(1)_{A}=U(1)_{\widetilde{\Phi}_{1}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$
The $T$-refined half-index in the A-twisted limit with (108) is given as
$\begin{split}I\!\!I\!(q,T,z)&=\frac{1}{(q)_{\infty}^{4}}\sum_{m\in\mathbb{Z}^{4}}q^{m_{1}(m_{2}+m_{3})}q^{m_{1}+m_{2}+m_{3}}z^{m_{4}-m_{2}}T^{m_{2}+m_{3}}\\ &\qquad\times\left(\prod_{i=1}^{4}(q^{1+m_{i}})_{\infty}\right)(q^{1+m_{1}}T)_{\infty}(q^{1+m_{2}+m_{3}-m_{4}})_{\infty}\\ &=(q)_{\infty}(qT)_{\infty}\sum_{m\in\mathbb{Z}^{4}}\frac{q^{m_{1}(m_{2}+m_{3})}q^{m_{1}+m_{2}+m_{3}}z^{m_{4}-m_{2}}T^{m_{2}+m_{3}}}{\left(\prod_{i=1}^{4}(q)_{m_{i}}\right)(qT)_{m_{1}}(q)_{m_{2}+m_{3}-m_{4}}}\ ,\end{split}$ (111)
@@
This agrees with the previous results.